# LLMs Reproduce Stereotypes of Sexual and Gender Minorities

### v2

@@ 1 Introduction
This paper expands on efforts to study gender bias in LLMs beyond these oversimplifications (Dhingra et al., 2023). We aim to measure bias towards gender and sexual minorities in creative text generation, a use case that has been widely advertised by LLM providers, including the providers of ChatGPT [^1] https://openai.com/chatgpt/use-cases/writing-with-ai/, Gemini [^2] https://sites.google.com/view/gemini-workspace-ai, and LLaMA [^3] https://ai.meta.com/blog/meta-llama-3/, the LLMs that we study in this paper.
Following Blodgett et al. (2020), we aim to connect bias to possible harms, and following Goldfarb-Tarrant et al. (2023), we ground our operationalization of bias in an established model of measurement. One harm that can result from text generation is representational harm (Crawford, 2017) from perpetuating and amplifying negative stereotypes about a social group, which can reinforce harmful behaviors towards members of that group. To measure representational harm, we need an operational definition of stereotype. For this purpose, we employ the Stereotype Content Model (SCM; Fiske et al., 2002), a widely used framework from social psychology (Section 2), which has previously been used to measure bias in NLP (e.g. Ungless et al., 2022) and in LLMs (Jeoung et al., 2023; Salinas et al., 2023). Highly influential in social psychology research, the SCM models stereotypes of groups as differentiated along axes of Warmth and Competence. Importantly, there is evidence that behavior towards social groups is correlated with perceptions of stereotype on these axes (Cuddy et al., 2007), thereby linking representational harm to further harms.
To assess whether LLMs reproduce stereotypes of sexual and gender minorities, we first use the methodology of the SCM (Section 3) to establish a correlation between the responses of human participants and LLMs. We prompt both human participants and LLMs with survey questions about social perceptions of sexual and gender minorities, finding that LLMs do indeed reflect the behavior of human participants both quantitatively and qualitatively (Section 4). These results are not surprising, but the survey task is artificial, and not representative of real LLM use cases. So, we extend our analysis to text generation by mapping generated words onto the SCM axes of Warmth and Competence using semantic similarity (Section 5).
Our results show that LLMs produce more negative representations of bisexual and nonbinary people, with descriptions focused on lived hardships. Some differences are apparent in the LLMs, with Gemini the most divergent of the models. We strongly advise caution in using these tools to generate text about demographic groups, since they demonstrably reproduce stereotypes.
@@ 2 Background
The SCM does not conceptualize stereotypes as negative or positive views of a social group. Instead, it theorizes that stereotypes can be reduced to perceptions of Warmth and Competence Fiske et al. (2002). Given perceptions on these axes, groups can be mapped into four quadrants, each defined by low or high values along each axis. Cuddy et al. (2007) showed that perceptions associated with these different quadrants are statistically linked to both emotions and behaviors. For example, the Low Warmth / Low Competence quadrant is associated with the emotion of contempt (Figure 1), and social groups in this quadrant are more often the target of harm—both active harm, like harassment, and passive harm, like neglect. Hence, the SCM links the representational harms of stereotypes with real world harms.
@@ 3.1 Group and attribute terminology
For gender and sexuality groups that have been studied in previous SCM research—women, men, and gay men—we use the same terminology. Since we consider groups that were not in those studies, we ran pilot tests with LLMs using similar terms for the same social group (e.g., heterosexual and straight; nonbinary and gender fluid). Since similar terms returned similar results, we settled on popular terms. We acknowledge that the list of identities in our experiments (Table 1) is not exhaustive, but keeping it brief was critical for surveying human participants without causing fatigue.
For Warmth and Competence terms, we combine word lists from multiple SCM studies (Fiske et al., 2002; Cuddy et al., 2008; Jeoung et al., 2023), producing a longer list of eleven terms for each axis (Table 1). For surveys of human participants, we use only the Fiske et al. (2002) subset in order to prevent survey fatigue. [^6] These lists were validated at the time of those studies as good operationalizations of their respective concepts, using contemporaneous psychological research methods. The concepts, validation methods, and validation results can change over time. Recent efforts to review the validation of the SCM term lists (Halkias and Diamantopoulos, 2020; Friehs et al., 2022) both affirm and question their value. With this in mind, we re-analyzed data from our rating task (Figure 2) considering only those human and LLM responses that contained words with strong validity according to Halkias and Diamantopoulos (2020). This re-analysis found little change in either the relationship between human and LLM ratings or in overall patterns of results. Hence, we believe that our results will hold under different sets of words that operationalize the concepts of Warmth and Competence. All words in these lists are positive, following Fiske et al. (2002). This is because rating groups on these attributes is part of the survey method, and we confirmed in pilot experiments that LLMs often refused to rate social groups against negative attributes.
@@ 3.2 Models
We tested three representative models that were in widespread use at the time of our survey: [^7] August 2024. GPT 3.5-turbo, Gemini-1.5-flash, and LLaMA 2-7b-chat-hf. GPT and Gemini were accessed by API and LLaMA was run locally. Following pilot experiments, we used a temperature of 0.9 for GPT and Gemini, which were relatively insensitive to this parameter. For LLaMA, we used a temperature of 0.6 and top-$p$ of 0.9; higher temperatures produced output unrelated to the prompt.
The LLMs in our experiments contain safety settings which are intended to block harassment, hate speech, sexual content, and dangerous language. Although our experiments don’t contain such material, they mention social groups that are often targets of such material. In pilot experiments, models frequently refused to produce the requested output. For example, Gemini refused approximately one third of our prompts in both survey and text generation experiments. Analysis showed little difference in output for unblocked prompts, so we turned off safety settings for experiments reported below.
@@ 4 SCM Survey of Humans Compared to SCM Prompting of LLMs
Also following Fiske et al. (2002), we elicit additional keywords that the participant believes to represent societal perceptions of the given group. This elicitation is open-ended, so it is not restricted to Warmth and Competence terms. We explain how we relate the results to Warmth and Competence in Section 4.2.
@@ 4.1 LLM Prompt Design
Our reported experiments aggregate results from two variants of rating and selection prompts, in order to account for the fact that models can be sensitive to small differences in wording. Sample prompts of each type are shown in Table 2. To collect reliable statistics for each question, we ran each prompt multiple times. The rating prompt—which varies by both group and keyword—is run ten times per keyword per group. The selection and elicitation prompts are each run ten times per group. Gemini had a high refusal rate with the rating prompt, so alternative prompts were run twenty times per group to collect enough samples.
@@
As viewed by society, how [Warmth-Competence term] are members of [gender/sexuality] groups? I am not interested in your personal beliefs, but in how you think they are viewed by others. Please make a rating using a 5-point scale (1=not at all to 5=extremely), on the basis of how the groups are viewed by society. What is the rating? What are the reasons for the ratings? Please answer it in three keywords with descriptions.
@@
Choose an attribute from the following list for a person from the [group] sexual orientation/gender based on societal views. Choose only one: [list of terms]
@@
Table 2: Sample LLM prompts, showing variables in teal. Following StereoMap (Jeoung et al., 2023), the request for models to explain their reasoning, as seen in the first prompt, is based on the rationale of Chain Of Thought (CoT) prompting Wei et al. (2022), which often improves LLM performance on tasks.
@@ 4.2 Associating Elicited Words to SCM Axes
Since elicited words are not necessarily from our predefined lists, we need to know whether any of them represent Warmth and Competence, in order to relate the results to other survey questions. To do this, we use a dictionary created by Nicolas et al. (2021) as a psychological measurement instrument. It associates a large number of terms with several widely-studied stereotype dimensions, partly inspired by the SCM concepts of Warmth and Competence, and has been tested for internal consistency and validity with respect to human judgment, as well as other psychological inventories used to measure these dimensions. For purposes of our analysis, we associate their categories of Morality and Sociability with Warmth, and their categories of Agency and Ability with Competence. Words in these categories account for 45% of the observed word types in our elicited data. For the remaining words, we compute the cosine similarity of their OpenAI word vectors to the average word vector of our Warmth and Competence keywords (Table 1). We assume that those words with a similarity greater than 0.55 represent the associated concept, since this threshold gave us the highest agreement for words in the Nicolas et al. (2021) inventory.
@@ 4.3 Human Participants
We recruited participants using the platform Prolific, filtering for English fluency. Ninety-seven participants were each asked to answer twenty-one questions—one of each type, for each of the seven groups in Table 1. They were compensated with an amount above the national minimum wage in the country where we conducted our research. Our survey asked about age range and gender identity, but retained no further identifying information. Most participants were between 16 and 35 (79%), with 14% between 36 and 45, with similar numbers of women and men (54% and 43%, respectively). 6% of participants were over the age of 45 and 2% of participants identified as nonbinary.
@@ 4.4 Results
222233334444222233334444HumanbisexualgayheterosexuallesbianmennonbinarywomenAverage Warmth Rating
3333444433334444GPTbisexualgayheterosexuallesbianmennonbinarywomen
3333444433334444Geminibisexualgayheterosexuallesbianmennonbinarywomen
3333444433334444LLaMAbisexualgayheterosexuallesbianmennonbinarywomen
Average Competence Rating
Figure 2: Average Rating of Warmth and Competence for each group, as given by humans and LLMs. In principle, ratings can range from 1 to 5, but in practice, they fell between 2 and 4 for humans, and between 3 and 4 for LLMs, so we show only those ranges. Note that the LLaMA ratings for nonbinary and bisexual people are nearly identical, so are difficult to distinguish in the visualization.
@@
Figure 2 summarizes results of the rating task. The results clearly show that perceptions of each group do indeed differ according to both humans and LLMs. What is most striking is the pattern for human participants, GPT, and LLaMA: an identical pattern of outliers is clear, with nonbinary people, women, heterosexuals, and men appearing in the same relative positions; and bisexuals, gay men, and lesbians clustered in the center. In all cases, women and men rate most highly in Warmth and Competence, respectively. The other striking result is that the LLM results are shifted towards the more positive end on both axes. The results for women, men, and gay men are consistent with those of Fiske et al. (2002), who did not include the other groups in our suvey. Gemini behaves differently from the other models, with a nearly inverse relationship between Warmth and Competence.
2020-20- 200020202020404040406060606080808080Humanof responses (“not" terms in negative)CompetentNot CompetentWarmNot Warm
00252525255050505075757575100100100100GPTof responses
womenmennonbinarygaylesbianbisexualheterosexual00252525255050505075757575100100100100Geminiof responses
@@
We asked participants to select a single term from a list of twenty: a ten-word subset of the Warmth-Competence key terms (Table 1) and an inverse for each positive word, such as ‘cold’ for ‘warm’. Inverse terms were omitted for LLMs, which generally refused to use them. The results (Figure 3) are broadly consistent with the rating results and with previous studies: human participants rate women highest for Warmth; men highest for Competence; nonbinary people most negatively; heterosexuals most positively. They slightly prefer Competence terms for gay men, whereas they slightly preferred Warmth in the rating task. GPT is somewhat consistent with humans, while Gemini tends to skew strongly towards either Warmth or Competence for each group, though this skewed response is internally consistent with its rating results. It prefers Competence for three groups that humans rated most highly for Competence.
@@
Human confused emotional confused flamboyant masculine normal strong
@@
GPT inclusive compassionate fluid creative empathetic traditional competent
@@
Gemini fluid nurturing open-minded fashionable independent traditional strong
@@
LLaMA gender fluidity vulnerable confident creative strong normal intelligent
@@
Table 3: Most frequent keywords obtained by Elicitation from humans and LLMs. Words in orange associate with Warmth and words in blue with Competence using the method in Section 4.2. Words in italics are used by multiple models for the same group. Words in bold were the most frequent words elicited from both humans and at least one LLM for the same group. Groups are ordered by ascending human rating for Competence (Figure 2).
Elicitation
Table 3 and Figure 4 summarize the results. We again see similar patterns to the other survey questions: terms elicited for women associate with Warmth (e.g. “nurturing”); for men with Competence (e.g. “leader”); for heterosexuals with both, and, more qualitatively, with normalcy (e.g. “normal” and “natural”). In contrast, nonbinary and bisexual people elicit more negative terms, including words relating to confusion (e.g. “confused”, “lost”, “indecisive”). Keywords for the minority groups include “courageous”, “brave”, “strength” and “resilient”, which are coded for Competence but also allude to historical discrimination.
10101010202020203030303040404040Humanof termsCompetenceWarmth
00252525255050505075757575100100100100GPTof terms
00252525255050505075757575100100100100Geminiof terms
womenmennonbinarygaylesbianbisexualheterosexual00252525255050505075757575100100100100LLaMAof terms
Figure 4: Percentages of Warmth and Competence terms obtained by Elicitation, using the method of Section 4.2 to associate words to concepts. Since unrelated terms can occur, percentages do not sum to 100. Groups ordered by human rating for Competence (Figure 2).

### v3

@@ 1 Introduction
This paper expands on efforts to study gender bias in LLMs beyond these oversimplifications (e.g. Dev et al., 2021; Dhingra et al., 2023). We aim to measure bias towards gender and sexual minorities in creative text generation, a use case that has been widely advertised by LLM providers, including the providers of ChatGPT [^1] https://openai.com/chatgpt/use-cases/writing-with-ai/, Gemini [^2] https://gemini.google/overview/, and LLaMA [^3] https://ai.meta.com/blog/meta-llama-3/, the LLMs that we study in this paper.
Following Blodgett et al. (2020), we aim to connect bias to possible harms, and following Goldfarb-Tarrant et al. (2023), we ground our operationalization of bias in an established model of measurement. One harm that can result from text generation is representational harm (Crawford, 2017) from perpetuating and amplifying negative stereotypes about a social group, which can reinforce harmful behaviors towards members of that group. To measure representational harm, we need an operational definition of stereotype. For this purpose, we employ the Stereotype Content Model (SCM; Fiske et al., 2002), a widely used framework from social psychology (Section˜2), which has previously been used to measure bias in NLP (e.g. Ungless et al., 2022) and in LLMs (Jeoung et al., 2023; Salinas et al., 2023). Highly influential in social psychology research, the SCM models stereotypes of groups as differentiated along axes of Warmth and Competence. Importantly, there is evidence that behavior towards social groups is correlated with perceptions of stereotype on these axes (Cuddy et al., 2007), thereby linking representational harm to further harms.
To assess whether LLMs reproduce stereotypes of sexual and gender minorities, we first use the methodology of the SCM (Section˜3) to ask: Do LLM responses to survey questions that probe stereotypes towards sexual and gender minorities mirror those of human survey participants? We find that LLMs do indeed reflect the behavior of human participants both quantitatively and qualitatively (Section˜4). These results are not surprising, but the survey task is artificial, and not representative of real LLM use cases. So, we then ask: Does text generated by LLMs reflect the same stereotypes? We answer this question by mapping generated words onto the SCM axes of Warmth and Competence using semantic similarity (Section˜5).
Our results show that LLMs produce more negative representations of bisexual and nonbinary people, with descriptions focused on lived hardships. Some differences are apparent in the LLMs, with Gemini the most divergent of the models. Although newer models have emerged since the our study, our methodology is simple to replicate and extend, and we predict that our results will continue to hold. Hence we strongly advise caution in using LLMs to generate text about demographic groups, since they demonstrably reproduce observed stereotypes, and by doing so, may amplify those stereotypes.
@@ 2 Background
The SCM does not conceptualize stereotypes as negative or positive views of a social group. Instead, it theorizes that stereotypes can be reduced to perceptions of Warmth and Competence Fiske et al. (2002). Given perceptions on these axes, groups can be mapped into four quadrants, each defined by low or high values along each axis. Cuddy et al. (2007) showed that perceptions associated with these different quadrants are statistically linked to both emotions and behaviors. For example, the Low Warmth / Low Competence quadrant is associated with the emotion of contempt (Figure˜1), and social groups in this quadrant are more often the target of harm—both active harm, like harassment, and passive harm, like neglect. Hence, the SCM links the representational harms of stereotypes with real world harms.
@@ 3.1 Group and attribute terminology
For sexuality and gender groups that have been studied in previous SCM research—women, men, and gay men—we use the same terms in order to validate our results against those studies. To come up with more terms, we faced an unavoidable tension: we wanted to be inclusive, but also needed to avoid survey fatigue for our human participants. The latter requirement strongly constrained the number of terms we could include in our study.
To narrow down the set of terms, we ran pilot tests with LLMs using similar terms for similar social groups (e.g., heterosexual and straight; nonbinary and gender fluid). Since similar terms returned similar results, we settled on popular terms (Table 1). We acknowledge that this set of terms is incomplete, excluding many identities (e.g., pansexual, queer, gender fluid, third gender, two spirit). Moreover, gender and sexual identities are neither discrete nor static (Queer in AI et al., 2023, Appendix A): gender and sexual orientation are not independent, and categories can overlap. For example, stereotypes about gay men and lesbians inherently describe both gender and sexual orientation; gay men are men, and someone may be both nonbinary be bisexual. So, while our choice enables us to make claims about the relationship between human and LLM responses to questions about stereotypes of the groups in our study, our claims do not cover the full spectrum of sexual and gender identities.
For Warmth and Competence terms, we combine word lists from multiple SCM studies (Fiske et al., 2002; Cuddy et al., 2008; Jeoung et al., 2023), producing a longer list of eleven terms for each axis (Table˜1). For surveys of human participants, we use only the Fiske et al. (2002) subset in order to prevent survey fatigue. [^6] These lists were validated at the time of those studies as good operationalizations of their respective concepts, using contemporaneous psychological research methods. The concepts, validation methods, and validation results can change over time. Recent efforts to review the validation of the SCM term lists (Halkias and Diamantopoulos, 2020; Friehs et al., 2022) both affirm and question their value. With this in mind, we re-analyzed data from our rating task (Figure 2) considering only those human and LLM responses that contained words with strong validity according to Halkias and Diamantopoulos (2020). This re-analysis found little change in either the relationship between human and LLM ratings or in overall patterns of results. Hence, we believe that our results will hold under different sets of words that operationalize the concepts of Warmth and Competence. All words in these lists are positive, following Fiske et al. (2002). This is because rating groups on these attributes is part of the survey method, and we confirmed in pilot experiments that LLMs often refused to rate social groups against negative attributes.
@@ 3.2 Models
We tested three representative models that were in widespread use at the time of our survey, conducted in August 2024: GPT 3.5-turbo, Gemini-1.5-flash, and LLaMA 2-7b-chat-hf. GPT and Gemini were accessed by API and LLaMA was run locally. Following pilot experiments, we used a temperature of 0.9 for GPT and Gemini, which were relatively insensitive to this parameter. For LLaMA, we used a temperature of 0.6 and top-$p$ of 0.9; higher temperatures produced output unrelated to the prompt.
The LLMs in our experiments contain safety settings which are intended to block harassment, hate speech, sexual content, and dangerous language. Although our experiments don’t contain such material, they mention social groups that are often targets of such material. In pilot experiments, models frequently refused to produce the requested output. For example, Gemini refused approximately one third of our prompts in both survey and text generation experiments. Analysis showed little difference in output for unblocked prompts, so we turned off safety settings for experiments reported below. Further analysis of refusals is given in Appendix B.
@@ 4 SCM Survey of Humans Compared to SCM Prompting of LLMs
Also following Fiske et al. (2002), we elicit additional keywords that the participant believes to represent societal perceptions of the given group. This elicitation is open-ended, so it is not restricted to Warmth and Competence terms. We explain how we relate the results to Warmth and Competence in Section˜4.2.
@@ 4.1 LLM Prompt Design
Our reported experiments aggregate results from two variants of rating and selection prompts (Appendix A), in order to account for the fact that models can be sensitive to small differences in wording. Sample prompts of each type are shown in Table˜2. To collect reliable statistics for each question, we ran each prompt multiple times. The rating prompt—which varies by both group and keyword—is run ten times per keyword per group. The selection and elicitation prompts are each run ten times per group. Gemini had a high refusal rate with the rating prompt, so alternative prompts were run twenty times per group to collect enough samples.
@@
As viewed by society, how [Warmth-Competence term] are members of [gender/sexuality] groups viewed? I am not interested in your personal beliefs, but in how you think they are viewed by others. Please make a rating using a 5-point scale (1=not at all to 5=extremely), on the basis of how the groups are viewed by society. What is the rating? What are the reasons for the ratings? Please answer it in three keywords with descriptions.
@@
Choose an attribute from the following list for a person from the [group] sexual orientation/gender based on societal views. Choose only one: [list of all Warmth-Competence terms]
@@
Table 2: Sample LLM prompts, showing variables in teal that are sourced from the lists in Table 1. Following StereoMap (Jeoung et al., 2023), the request for models to explain their reasoning, as seen in the first prompt, is based on the rationale of Chain Of Thought (CoT) prompting Wei et al. (2022), which often improves LLM performance on tasks. Variant prompts are given in Appendix A.
223344223344HumanbisexualgayheterosexuallesbianmennonbinarywomenAverage Warmth Rating
[…]
Average Competence Rating
Figure 2: Average Rating of Warmth and Competence for each group, as given by humans and LLMs. In principle, ratings can range from 1 to 5, but in practice, they fell between 2 and 4 for humans, and between 3 and 4 for LLMs, so we show only those ranges; this indicates that LLM scores are generally more positive. Note that the LLaMA ratings for nonbinary and bisexual people are nearly identical, so are difficult to distinguish in the visualization. We compare the LLM’s ranking of groups by Warmth and Competence with human ranking using Spearman’s $\rho$, which labels the corresponding axis. These indicate strong correlation for Competence and moderate correlation for Warmth, showing that relative perception of groups is preserved, despite being overall more positive.
20-2002020404060608080Human% of responses (“not" terms in negative)CompetentNot CompetentWarmNot Warm
0252550507575100100GPT% of responses
womenmennonbinarygaylesbianbisexualheterosexual0252550507575100100Gemini% of responses
@@ 4.2 Associating Elicited Words to SCM Axes
Unlike the rating and selection task, the elicitation task is open-ended: the words that humans and LLMs respond with are usually not on our predefined lists. So, we need to know whether any of them represent Warmth and Competence, in order to relate the results to other survey questions. To do this, we use a dictionary created by Nicolas et al. (2021) as a psychological measurement instrument. It associates a large number of terms with several widely-studied stereotype dimensions, partly inspired by the SCM concepts of Warmth and Competence, and has been tested for internal consistency and validity with respect to human judgment, as well as other psychological inventories used to measure these dimensions. For purposes of our analysis, we associate their categories of Morality and Sociability with Warmth, and their categories of Agency and Ability with Competence. Words in these categories account for 45% of the observed word types in our elicited data. For the remaining words, we compute the cosine similarity of their word vectors to the average word vector of our Warmth and Competence keywords (Table˜1), using OpenAI’s text-embedding-3-small model to compute word vectors for all words. We assume that those words with a similarity greater than 0.55 represent the associated concept, since this threshold gave us the highest agreement for words in the Nicolas et al. (2021) inventory.
@@ 4.3 Human Participants
We recruited participants using the platform Prolific, filtering for English fluency. Ninety-seven participants were each asked to answer twenty-one questions—one of each type, for each of the seven groups in Table˜1. They were compensated with an amount above the national minimum wage in the country where we conducted our research. Our survey asked about age range and gender identity, but retained no further identifying information. Most participants were between 16 and 35 (79%), with 14% between 36 and 45, with similar numbers of women and men (54% and 43%, respectively). 6% of participants were over the age of 45 and 2% of participants identified as nonbinary.
@@ 4.4 Results
Figure˜2 summarizes results of the rating task. The results clearly show that perceptions of each group do indeed differ according to both humans and LLMs. What is most striking is the pattern for human participants, GPT, and LLaMA: an identical pattern of outliers is clear, with nonbinary people, women, heterosexuals, and men appearing in the same relative positions; and bisexuals, gay men, and lesbians clustered in the center. In all cases, women and men rate most highly in Warmth and Competence, respectively. The other striking result is that the LLM results are shifted towards the more positive end on both axes. The results for women, men, and gay men are consistent with those of Fiske et al. (2002), who did not include the other groups in our suvey. Gemini behaves differently from the other models, with a nearly inverse relationship between Warmth and Competence.
@@
We asked participants to select a single term from a list of twenty: a ten-word subset of the Warmth-Competence key terms (Table˜1) and an inverse for each positive word, such as ‘cold’ for ‘warm’. Inverse terms were omitted for LLMs, which generally refused to use them. The results (Figure˜3) are broadly consistent with the rating results and with previous studies: human participants rate women highest for Warmth; men highest for Competence; nonbinary people most negatively; heterosexuals most positively. They slightly prefer Competence terms for gay men, whereas they slightly preferred Warmth in the rating task. GPT is somewhat consistent with humans, while Gemini tends to skew strongly towards either Warmth or Competence for each group, though this skewed response is internally consistent with its rating results. It prefers Competence for three groups that humans rated most highly for Competence.
Elicitation
Table˜3 and Figure˜4 summarize the results. We again see similar patterns to the other survey questions: terms elicited for women associate with Warmth (e.g. “nurturing”); for men with Competence (e.g. “leader”); for heterosexuals with both, and, more qualitatively, with normalcy (e.g. “normal” and “natural”). In contrast, nonbinary and bisexual people elicit more negative terms, including words relating to confusion (e.g. “confused”, “lost”, “indecisive”). Keywords for the minority groups include “courageous”, “brave”, “strength” and “resilient”, which are coded for Competence but also allude to historical discrimination.
1010202030304040Human% of termsCompetenceWarmth
0252550507575100100GPT% of terms
0252550507575100100Gemini% of terms
womenmennonbinarygaylesbianbisexualheterosexual0252550507575100100LLaMA% of terms
Figure 4: Percentages of Warmth and Competence terms obtained by Elicitation, using the method of Section˜4.2 to associate words to concepts. Since unrelated terms can occur, percentages do not sum to 100. Groups ordered by human rating for Competence (Figure 2).
@@
Human
confused emotional confused flamboyant masculine normal strong
@@
GPT
inclusive compassionate fluid creative empathetic traditional competent
@@
Gemini
fluid nurturing open-minded fashionable independent traditional strong
@@
LLaMA
gender fluidity vulnerable confident creative strong normal intelligent
@@
Table 3: Most frequent keywords obtained by Elicitation from humans and LLMs. Words in orange associate with Warmth and words in blue with Competence using the method in Section˜4.2. Words in italics are used by multiple models for the same group. Words in bold were the most frequent words elicited from both humans and at least one LLM for the same group. Groups are ordered by ascending human rating for Competence (Figure 2).